# Directional data analysis using the spherical Cauchy and the Poisson kernel-based distribution

### v4

@@ 3.3 Regression analysis
This procedure was iterated for various combinations of sample sizes and dimensionalities, with regression coefficients being estimated using both the SC and PKB regression models. The entire process was repeated 1,000 times, and the average fit of the two models, as measured by the quantity $\sum_{i=1}^{n}{\bf y}_{i}^{\top}\hat{\bf y}_{i}/n$ is presented in Table 4 The fit takes values from 0 up to 1, where higher values indicate better fit. The estimated fits of the regression models are nearly nearly identical.
Table 4: Estimated fit of each regression model, computed as $\sum_{i=1}^{n}{\bf y}_{i}^{\top}\hat{\bf y}_{i}/n$ when the data are generated from the SC and the PKB distribution.
@@ 3.4 Discriminant analysis
Table 5: Estimated percentage of correct classification when the data are generated from the SC and the PKB distribution, and the concentration parameter was equal to $\rho=0.5$. The angle between the location parameters of the two groups is denoted by $\theta$.
@@ 3.5 Model based clustering
Table 6: Estimate of the absolute error $|\hat{K}-K|$ where $\hat{K}$ and $K$ denote the estimated and true number of clusters, respectively. We considered various combinations of dimensionalities ($d$), true number of clusters ($K$) and sample size $n$. In all cases the data are generated from the SC and the PKB distribution.
@@
$K$ SC PKBD SC PKBD SC PKBD SC PKBD SC PKBD SC PKBD SC PKBD SC PKBD
@@
Table 7: Average ARI for various combinations of dimensionalities ($d$), true number of clusters ($K$) and sample size $n$. In all cases the data are generated from the SC and the PKB distribution.
@@ 4.2 Regression analysis
Data concerning crop productivity in the Greek NUTS II region of Thessaly during the 2017–2018 cropping year were provided by the Greek Ministry of Agriculture, specifically through the farm accountancy data network (FADN). The dataset pertains to a sample of 487 farms and initially included 20 crops. However, after aggregation, the number of crops was reduced to 10666A larger version of this dataset was employed in (Mattas et al.,, 2024).. For each of the 487 farms, both the cultivated area and the production for each of the 10 crops are recorded. The objective of this study is to relate the composition of production (simplicial response, $\mathbf{Y}$) to the composition of cultivated area (simplicial predictor, $\mathbf{X}$). Consequently, the data were scaled to sum to unity [^7] The raw data cannot be distributed due to disclosure restrictions.. A square root transformation was applied to the compositional data (both the response and predictor variables), mapping them onto a 10-dimensional sphere.
@@ 4.4 Model based clustering
We utilize the Wireless Indoor Localization dataset again, on the grounds that it is a hard case, as pointed out by Golzy and Markatou, (2020) as well. We searched for the optimal number of components (clusters) using the BIC (Schwarz,, 1978) and ICL Biernacki et al., (2000) criteria.
@@ 5 Conclusions
We investigated two recently proposed distributions, namely the spherical Cauchy (SC) and the Poisson kernel-based (PKB) distributions. We introduced an alternative, hybrid method for estimating the parameters of both distributions. For the SC distribution, we proposed a new parameterization that facilitates the application of the Newton-Raphson algorithm. Additionally, we re-parameterized the density function of the PKB distribution in a similar manner to the SC distribution. However, this re-parameterization did not prove useful for maximum likelihood estimation (MLE) of the PKB parameters. Nevertheless, the new parameterizations enabled the implementation of regression models by circumventing the (hyper-)spherical constraint, thus allowing for anisotropic errors similar to those found in the ESAG distribution (Paine et al.,, 2020). The advantage of the re-parameterized SC distribution is that the Newton-Raphson method could also be applied in the regression context, leading to computationally efficient estimation of the regression parameters. The hybrid estimation method for both distributions further enabled hypothesis testing of location directions between two populations using the log-likelihood ratio test, without the assumption of equal concentration parameters. Lastly, we explored classification as well as model-based clustering under a maximum likelihood framework using either distribution.
@@
$=$ $\frac{d-1}{2}\left[\log{\frac{1+\rho^{2}-2{\bf y}^{\top}{\bf m}}{1-\rho^{2}}}+\log{\left(1-\rho^{2}\right)}\right]$
$=$ $\frac{d-1}{2}\log{\left(1+\rho^{2}-2{\bf y}^{\top}{\bf m}\right)}.$
A2 Derivatives of Eq. (5)
${\bf J}$ $=$ $\frac{\partial\ell_{SC}}{\partial\boldsymbol{\mu}}=-d\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-{\bf y}_{i}}{\sqrt{\gamma^{2}+1}-\alpha_{i}}$
${\bf H}$ $=$ $\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\mu}\partial\boldsymbol{\mu}^{\top}}=-d\sum_{i=1}^{n}\dfrac{\frac{{\bf I}_{d+1}\sqrt{\gamma^{2}+1}-\frac{\boldsymbol{\mu}\boldsymbol{\mu}^{\top}}{\sqrt{\gamma^{2}+1}}}{\gamma^{2}+1}\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)-\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-{\bf y}_{i}\right)\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-{\bf y}_{i}\right)^{\top}}{\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)^{2}}.$
A3 Derivatives of Eq. (8)
${\bf J}=\frac{\partial\ell_{PKB}}{\partial\boldsymbol{\mu}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-{\bf y}_{i}}{\sqrt{\gamma^{2}+1}-\alpha_{i}}-n\frac{d-1}{2}\left(\dfrac{\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}}{\sqrt{\gamma^{2}+1}-1}-\frac{2\boldsymbol{\mu}}{\gamma^{2}}\right)$
${\bf H}=\frac{\partial^{2}\ell_{PKB}}{\partial\boldsymbol{\mu}\partial\boldsymbol{\mu}^{\top}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\dfrac{\frac{{\bf I}_{d+1}\sqrt{\gamma^{2}+1}-\frac{\boldsymbol{\mu}\boldsymbol{\mu}^{\top}}{\sqrt{\gamma^{2}+1}}}{\gamma^{2}+1}\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)-\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-{\bf y}_{i}\right)\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-{\bf y}_{i}\right)^{\top}}{\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)^{2}}$
@@
A4 Derivatives of Eq. (11)
$\frac{\partial\ell_{SC}}{\partial\boldsymbol{\beta}_{k}}$ $=$ $-d\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}}$
$\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\beta}_{k}\partial\boldsymbol{\beta}_{l}^{\top}}$ $=$ $\left\{\begin{array}[]{cc}-d\sum_{i=1}^{n}\frac{\frac{{\bf x}_{i}{\bf x}_{i}^{\top}\sqrt{\gamma_{i}^{2}+1}-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{ik}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}},&\text{if}\ k=l\\ -d\sum_{i=1}^{n}\frac{\frac{-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{il}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{il}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{il}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}},&\text{if}\ k\neq l\\ \end{array}\right\}$
A5 Derivatives of Eq. (12)
@@
$\frac{\partial\ell_{PKB}}{\partial\boldsymbol{\beta}_{k}}=-\frac{d+1}{2}\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}}-\frac{d-1}{2}\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}}}{\sqrt{\gamma_{i}^{2}+1}-1}+\frac{d-1}{2}\sum_{i=1}^{n}\frac{2\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\gamma_{i}^{2}}.$
@@
$\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\beta}_{k}\partial\boldsymbol{\beta}_{k}^{\top}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\frac{\frac{{\bf x}_{i}{\bf x}_{i}^{\top}\sqrt{\gamma_{i}^{2}+1}-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{ik}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}}$
@@
$\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\beta}_{k}\partial\boldsymbol{\beta}_{l}^{\top}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\frac{\frac{-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{il}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{il}}{\sqrt{\gamma_{i}^{2}+1}}-{\bf y}_{il}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}}$

### v5

@@ 3.3 Regression analysis
This procedure was iterated for various combinations of sample sizes and dimensionalities, with regression coefficients being estimated using both the SC and PKB regression models. The entire process was repeated 1,000 times, and the average fit of the two models, as measured by the quantity $\sum_{i=1}^{n}\bm{y}_{i}^{\top}\hat{\bm{y}}_{i}/n$ is presented in Table 4. The fit takes values from 0 up to 1, where higher values indicate better fit. The estimated fits of the regression models are nearly nearly identical.
Table 4: Estimated fit of each regression model, computed as $\sum_{i=1}^{n}\bm{y}_{i}^{\top}\hat{\bm{y}}_{i}/n$ when the data are generated from the SC and the PKB distributions.
@@ 3.4 Discriminant analysis
Table 5: Estimated percentage of correct classification when the data are generated from the SC and the PKB distributions, and the concentration parameter was equal to $\rho=0.5$. The angle between the location parameters of the two groups is denoted by $\theta$.
@@ 3.5 Model based clustering
Table 6: Estimate of the absolute error $|\hat{K}-K|$ where $\hat{K}$ and $K$ denote the estimated and true number of clusters, respectively. We considered various combinations of dimensionalities ($d$), true number of clusters ($K$) and sample size $n$. In all cases the data are generated from the SC and the PKB distributions.
@@
$K$ SC PKB SC PKB SC PKB SC PKB SC PKB SC PKB SC PKB SC PKB
@@
Table 7: Average ARI for various combinations of dimensionalities ($d$), true number of clusters ($K$) and sample size $n$. In all cases the data are generated from the SC and the PKB distributions.
@@ 4.2 Regression analysis
Data concerning crop productivity in the Greek NUTS II region of Thessaly during the 2017–2018 cropping year were provided by the Greek Ministry of Agriculture, specifically through the farm accountancy data network (FADN). The dataset pertains to a sample of 487 farms and initially included 20 crops. However, after aggregation, the number of crops was reduced to 10666A larger version of this dataset was employed in Mattas et al., (2024).. For each of the 487 farms, both the cultivated area and the production for each of the 10 crops are recorded. The objective of this study is to relate the composition of production (simplicial response, $\mathbf{Y}$) to the composition of cultivated area (simplicial predictor, $\mathbf{X}$). Consequently, the data were scaled to sum to unity [^7] The raw data cannot be distributed due to disclosure restrictions.. A square root transformation was applied to the compositional data (both the response and predictor variables), mapping them onto a 10-dimensional sphere.
@@ 4.4 Model based clustering
We utilize the Wireless Indoor Localization dataset again, on the grounds that it is a hard case, as pointed out by Golzy and Markatou, (2020) as well. We searched for the optimal number of components (clusters) using the BIC (Schwarz,, 1978) and ICL (Biernacki et al.,, 2000) criteria.
@@ 5 Conclusions
We investigated two recently proposed distributions, namely the spherical Cauchy (SC) and the Poisson kernel-based (PKB) distributions. We introduced an alternative, hybrid method for estimating the parameters of both distributions. For the SC distribution, we proposed a new parameterization that facilitates the application of the Newton-Raphson algorithm. Additionally, we re-parameterized the density function of the PKB distribution in a similar manner to the SC distribution. However, this re-parameterization did not prove useful for maximum likelihood estimation (MLE) of the PKB parameters. Nevertheless, the new parameterizations enabled the implementation of regression models by circumventing the (hyper-)spherical constraint, thus allowing for heterosceadstic errors similar to those found in the ESAG distribution (Paine et al.,, 2020). The advantage of the re-parameterized SC distribution is that the Newton-Raphson method could also be applied in the regression context, leading to computationally efficient estimation of the regression parameters. The hybrid estimation method for both distributions further enabled hypothesis testing of location directions between two populations using the log-likelihood ratio test, without the assumption of equal concentration parameters. Lastly, we explored classification as well as model-based clustering under a maximum likelihood framework using either distribution.
@@
$=$ $\frac{d-1}{2}\left[\log{\frac{1+\rho^{2}-2\bm{y}^{\top}{\bf m}}{1-\rho^{2}}}+\log{\left(1-\rho^{2}\right)}\right]$
$=$ $\frac{d-1}{2}\log{\left(1+\rho^{2}-2\bm{y}^{\top}{\bf m}\right)}.$
A2 Derivatives of the log-likelihood of SC, Eq. (5)
${\bf J}$ $=$ $\frac{\partial\ell_{SC}}{\partial\boldsymbol{\mu}}=-d\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-\bm{y}_{i}}{\sqrt{\gamma^{2}+1}-\alpha_{i}}$
${\bf H}$ $=$ $\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\mu}\partial\boldsymbol{\mu}^{\top}}=-d\sum_{i=1}^{n}\dfrac{\frac{{\bf I}_{d+1}\sqrt{\gamma^{2}+1}-\frac{\boldsymbol{\mu}\boldsymbol{\mu}^{\top}}{\sqrt{\gamma^{2}+1}}}{\gamma^{2}+1}\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)-\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-\bm{y}_{i}\right)\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-\bm{y}_{i}\right)^{\top}}{\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)^{2}}.$
A3 Derivatives of the log-likelihood of PKB, Eq. (8)
${\bf J}=\frac{\partial\ell_{PKB}}{\partial\boldsymbol{\mu}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-\bm{y}_{i}}{\sqrt{\gamma^{2}+1}-\alpha_{i}}-n\frac{d-1}{2}\left(\dfrac{\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}}{\sqrt{\gamma^{2}+1}-1}-\frac{2\boldsymbol{\mu}}{\gamma^{2}}\right)$
${\bf H}=\frac{\partial^{2}\ell_{PKB}}{\partial\boldsymbol{\mu}\partial\boldsymbol{\mu}^{\top}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\dfrac{\frac{{\bf I}_{d+1}\sqrt{\gamma^{2}+1}-\frac{\boldsymbol{\mu}\boldsymbol{\mu}^{\top}}{\sqrt{\gamma^{2}+1}}}{\gamma^{2}+1}\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)-\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-\bm{y}_{i}\right)\left(\frac{\boldsymbol{\mu}}{\sqrt{\gamma^{2}+1}}-\bm{y}_{i}\right)^{\top}}{\left(\sqrt{\gamma^{2}+1}-\alpha_{i}\right)^{2}}$
@@
A4 Derivatives of the log-likelihood of SC with covariates, Eq. (11)
$\frac{\partial\ell_{SC}}{\partial\boldsymbol{\beta}_{k}}$ $=$ $-d\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}}$
$\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\beta}_{k}\partial\boldsymbol{\beta}_{l}^{\top}}$ $=$ $\left\{\begin{array}[]{cc}-d\sum_{i=1}^{n}\frac{\frac{{\bf x}_{i}{\bf x}_{i}^{\top}\sqrt{\gamma_{i}^{2}+1}-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{ik}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}},&\text{if}\ k=l\\ -d\sum_{i=1}^{n}\frac{\frac{-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{il}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{il}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{il}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}},&\text{if}\ k\neq l\\ \end{array}\right\}$
A5 Derivatives of the log-likelihood of PKB with covariates, Eq. (14)
@@
$\frac{\partial\ell_{PKB}}{\partial\boldsymbol{\beta}_{k}}=-\frac{d+1}{2}\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}}-\frac{d-1}{2}\sum_{i=1}^{n}\dfrac{\frac{\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\sqrt{\gamma_{i}^{2}+1}}}{\sqrt{\gamma_{i}^{2}+1}-1}+\frac{d-1}{2}\sum_{i=1}^{n}\frac{2\boldsymbol{\mu}_{ik}{\bf x}_{i}}{\gamma_{i}^{2}}.$
@@
$\frac{\partial^{2}\ell_{PKB}}{\partial\boldsymbol{\beta}_{k}\partial\boldsymbol{\beta}_{k}^{\top}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\frac{\frac{{\bf x}_{i}{\bf x}_{i}^{\top}\sqrt{\gamma_{i}^{2}+1}-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{ik}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}}$
@@
$\frac{\partial^{2}\ell_{PKB}}{\partial\boldsymbol{\beta}_{k}\partial\boldsymbol{\beta}_{l}^{\top}}$ $=$ $-\frac{d+1}{2}\sum_{i=1}^{n}\frac{\frac{-\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}\boldsymbol{\mu}_{il}{\bf x}_{i}^{\top}}{\sqrt{\gamma_{i}^{2}+1}}}{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{ik}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{ik}{\bf x}_{i}\right)\left(\frac{{\bf x}_{i}\boldsymbol{\mu}_{il}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{il}{\bf x}_{i}\right)^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}}$
@@
A6 Asymptotic normality of the maximum likelihood estimators for the regression models in Section 2.4
Under certain regularity conditions (see, e.g., van der Vaart,, 1998, Theorem 5.39), general theory implies that the asymptotic normality, specifically (12) and (15), hold for the maximum likelihood estimators of the regression models given in Section 2.4.
The observed Fisher information matrices, ${\cal I}_{SC}$ and ${\cal I}_{PKB}$, which appear in (12) and (15), respectively, can be calculated as follows. It is straightforward to see that ${\cal I}_{SC}$ takes the form
${\cal I}_{SC}$ $=-\frac{1}{n}\frac{\partial^{2}\ell_{SC}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}$
$=\frac{d}{n}\sum_{i=1}^{n}\frac{\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)-\frac{\partial}{\partial\boldsymbol{\beta}}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)\frac{\partial}{\partial\boldsymbol{\beta}}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{\top}}{(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i})^{2}}.$ (A6.2)
The components of the derivatives in (A6.2), required to calculate ${\cal I}_{SC}$, are provided in Appendix A4. Using these components, the derivatives in (A6.2) can be expressed as
$\frac{\partial}{\partial\boldsymbol{\beta}}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)=\frac{1}{\sqrt{\gamma_{i}^{2}+1}}\begin{pmatrix}(\boldsymbol{\beta}_{1}^{\top}\bm{x}_{i})\bm{x}_{i}\\ \vdots\\ (\boldsymbol{\beta}_{d+1}^{\top}\bm{x}_{i})\bm{x}_{i}\end{pmatrix}-\begin{pmatrix}y_{i1}\bm{x}_{i}\\ \vdots\\ y_{i,d+1}\bm{x}_{i}\end{pmatrix}=\left(\frac{\bm{\mu}_{i}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{i}\right)\otimes\bm{x}_{i},$
$\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)$ $=\frac{\partial}{\partial\boldsymbol{\beta}^{\top}}\left(\frac{\bm{\mu}_{i}}{\sqrt{\gamma_{i}^{2}+1}}-\bm{y}_{i}\right)\otimes\bm{x}_{i}$
$=\frac{\sqrt{\gamma_{i}^{2}+1}\frac{\partial}{\partial\boldsymbol{\beta}^{\top}}\left(\bm{\mu}_{i}\otimes\bm{x}_{i}\right)-\left(\bm{\mu}_{i}\otimes\bm{x}_{i}\right)\frac{\partial}{\partial\bm{\beta}^{\top}}\sqrt{\gamma_{i}^{2}+1}}{\gamma_{i}^{2}+1}$
$=\frac{\sqrt{\gamma_{i}^{2}+1}\,{\bf I}_{(d+1)p}\otimes\bm{x}_{i}^{\top}\otimes\bm{x}_{i}-(\bm{\mu}_{i}\otimes\bm{x}_{i})(\bm{\mu}_{i}\otimes\bm{x}_{i})^{\top}}{\gamma_{i}^{2}+1}.$
Substituting these results into (A6.2), we obtain the observed Fisher information in the form of (13).
The observed Fisher information ${\cal I}_{PKB}$ for the PKB-based regression model can be derived from (14) as
$\begin{split}{\cal I}_{PKB}=-\frac{1}{n}\frac{\partial^{2}\ell_{PKB}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}=&-\frac{d+1}{2n}\sum_{i=1}^{n}\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\log\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)\\ &-\frac{d-1}{2n}\sum_{i=1}^{n}\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\left\{\log\left(\sqrt{\gamma_{i}^{2}+1}-1\right)-\log\gamma_{i}^{2}\right\}.\end{split}$ (A6.3)
Note that the calculations for ${\cal I}_{SC}$ and the results in Appendix A5 imply
$\sum_{i=1}^{n}\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\log\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)=\frac{n}{d}\,{\cal I}_{SC},$
$\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\log\left(\sqrt{\gamma_{i}^{2}+1}-1\right)$ $=\frac{\left(\sqrt{\gamma_{i}^{2}+1}-1\right)\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\sqrt{\gamma_{i}^{2}+1}-\frac{\partial}{\partial\boldsymbol{\beta}}\sqrt{\gamma_{i}^{2}+1}\frac{\partial}{\partial\boldsymbol{\beta}^{\top}}\sqrt{\gamma_{i}^{2}+1}}{\left(\sqrt{\gamma_{i}^{2}+1}-1\right)^{2}}$
$=\frac{\left(\sqrt{\gamma_{i}^{2}+1}-1\right){\bf I}_{d+1}\otimes\bm{x}_{i}^{\top}\otimes\bm{x}_{i}-(\gamma_{i}^{2}+1)^{-1}(\bm{\mu}_{i}\otimes\bm{x}_{i})(\bm{\mu}_{i}\otimes\bm{x}_{i})^{\top}}{\left(\sqrt{\gamma_{i}^{2}+1}-1\right)^{2}},$
$\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\log\gamma_{i}^{2}$ $=\gamma_{i}^{-4}\left(\gamma_{i}^{2}\frac{\partial^{2}}{\partial\boldsymbol{\beta}\partial\boldsymbol{\beta}^{\top}}\gamma_{i}^{2}-\frac{\partial}{\partial\boldsymbol{\beta}}\gamma_{i}^{2}\frac{\partial}{\partial\boldsymbol{\beta}^{\top}}\gamma_{i}^{2}\right)$
$=\gamma_{i}^{-4}\left\{2\gamma_{i}^{2}{\bf I}_{d+1}\otimes\bm{x}_{i}^{\top}\otimes\bm{x}_{i}-4(\bm{\mu}_{i}\otimes\bm{x}_{i})(\bm{\mu}_{i}\otimes\bm{x}_{i})^{\top}\right\}.$
Substitution of these results into (A6.3) leads to the observed Fisher information expressed as (16).